### v2

# The Tension of Space as Dark Energy: A No Geometric Sequestering Theorem for Minimal Sequesters

Abstract
We model space as an elastic membrane and identify its uniform tension $T_{s}$ with the vacuum energy density of space. The central result is a no geometric sequestering theorem. In any minimal matter vacuum sequester framework where a global constraint cancels constant contributions from the matter sector by adjusting a global variable, a purely geometric unit operator remains intact. Concretely, a volume term of the form $S_{s}=-T_{s}\int\sqrt{-g}\,d^{4}x$ does not couple to the global variable that enforces the matter cancellation, and it survives unchanged in the local field equations. We establish this result within a covariant effective field theory using the background field method. The analysis shows that the coefficient $T_{s}$ is renormalized by graviton loops and in general by mixed matter–graviton loops and therefore runs with the renormalization scale $\mu$. The running defines an effective cosmological term $T_{s}(\mu)$ that is present even after the matter vacuum energy has been neutralized by the sequester constraint. In particular, we find that $T_{s}(\mu)$ sources $G_{\mu\nu}=-T_{s}(\mu)\,g_{\mu\nu}+\cdots$. The tension term $T_{s}(\mu)$ can then be understood as residual dark energy in the minimal sequestering framework.
1 Introduction
When Einstein wrote the field equations he and many contemporaries believed the universe was static. A closer inspection showed that with homogeneous nonzero energy density a static solution does not exist for the original equations. In 1917 he added a cosmological term $\Lambda g_{\mu\nu}$ [2] to allow a static closed universe, which leads to
$R_{\mu\nu}-\frac{1}{2}R\,g_{\mu\nu}+\Lambda g_{\mu\nu}=\frac{8\pi G}{c^{4}}T_{\mu\nu}$ (1.1)
With this term it was clear that a homogeneous and isotropic closed universe with nonzero energy density $\rho$ can be static if $\Lambda=1/a^{2}=4\pi G\rho/c^{2}$, where $a$ is the curvature radius. After Hubble’s discovery of cosmic expansion [3] and the theoretical developments of Friedmann [4] and Lemaître [5], Einstein abandoned the cosmological term, reportedly calling it his biggest blunder.
Late in the 1990s the situation changed. Two supernova projects, the Supernova Cosmology Project [7] and the High-Z Supernova Search Team [6], measured Type Ia supernovae as standardizable candles [^1] Type Ia supernovae were used as standard candles due to their remarkably uniform light curves. and found them to be dimmer than expected in a decelerating model. The inferred expansion is accelerating. Independent evidence from the Cosmic Microwave Background anisotropies by DASI [8], WMAP [9], and Planck [10], and from Baryon Acoustic Oscillations [11, 12], supports this conclusion. For a classic review see [13].
Vacuum Energy and the Cosmological Term
The cosmological term invites an interpretation as vacuum energy. The vacuum should look the same to all inertial observers, so any stress it contributes must be Lorentz invariant in locally inertial frames. The only symmetric rank two tensor with this property is $\eta_{\mu\nu}$, which implies,
$T^{(\mathrm{vac})}_{\mu\nu}=-\rho_{\mathrm{vac}}\,c^{2}\,\eta_{\mu\nu}$ (1.2)
In arbitrary coordinates this becomes,
$T^{(\mathrm{vac})}_{\mu\nu}=-\rho_{\mathrm{vac}}\,c^{2}\,g_{\mu\nu}$ (1.3)
with $\rho_{\mathrm{vac}}\,c^{2}$ constant in spacetime (to preserve Lorentz invariance [14]).
The same form follows from a perfect fluid with equation of state $p=\omega\rho c^{2}$. The fluid stress tensor is [14],
$T_{\mu\nu}=\left(\rho+\frac{p}{c^{2}}\right)\,u_{\mu}\,u_{\nu}+p\,g_{\mu\nu}$ (1.4)
Using a barotropic equation of state $\omega=-1$, we have $p=-\rho_{\mathrm{vac}}\,c^{2}$ and therefore we obtain for the vacuum,
$T^{(\mathrm{vac})}_{\mu\nu}=\left(\rho_{vac}-\rho_{vac}\right)\,u_{\mu}\,u_{\nu}-{\rho_{vac}\ c^{2}}g_{\mu\nu}=-\rho_{\mathrm{vac}}\,c^{2}\,g_{\mu\nu}$ (1.5)
Thus a vacuum with constant energy density acts gravitationally like a perfect fluid with $\omega=-1$. It is natural to split the total stress tensor into matter plus radiation and vacuum parts,
$T_{\mu\nu}=T^{(M+R)}_{\mu\nu}+T^{(\mathrm{vac})}_{\mu\nu}$ (1.6)
Inserting this into Einstein’s original equations without $\Lambda$, we get,
$R_{\mu\nu}-\frac{1}{2}R\,g_{\mu\nu}=\frac{8\pi G}{c^{4}}\left(T^{(M+R)}_{\mu\nu}+T^{(\mathrm{vac})}_{\mu\nu}\right)$ (1.7)
and using $T^{(\mathrm{vac})}_{\mu\nu}=-\rho_{\mathrm{vac}}\,c^{2}\,g_{\mu\nu}$, we obtain,
$R_{\mu\nu}-\frac{1}{2}R\,g_{\mu\nu}=\frac{8\pi G}{c^{4}}T^{(M+R)}_{\mu\nu}-\frac{8\pi G\rho_{\mathrm{vac}}}{c^{2}}g_{\mu\nu}$ (1.8)
Defining $\Lambda\equiv 8\pi G\rho_{\mathrm{vac}}/c^{2}$, one recovers the equations with a cosmological term,
$R_{\mu\nu}-\frac{1}{2}R\,g_{\mu\nu}+\Lambda g_{\mu\nu}=\frac{8\pi G}{c^{4}}T^{(M+R)}_{\mu\nu}$ (1.9)
Hence the cosmological term is equivalent to vacuum energy in the field equations. Einstein introduced $\Lambda$ for a static model, but the same term arises from the presence of vacuum energy even without that motivation.
Dark Energy and the Cosmological Constant Problem
The vacuum contribution that drives late time acceleration is often called dark energy. It is dark because it does not radiate or absorb in a way we can directly detect, and its evidence comes from dynamics such as the acceleration rate. The measured mean energy density today is about $5.249\times 10^{-10}\,\mathrm{J\,m^{-3}}$. A naive quantum field theory estimate of the zero point energy gives a value larger by about $122$ orders of magnitude. This huge discrepancy is the cosmological constant problem. The question it raises is simple to state but hard to answer. What sets the tiny value of the vacuum energy that gravitates on cosmic scales, if the quantum contributions of known fields appear so large in simple estimates?
Many ideas have been explored. Representative proposals include quintessence [15, 16], k-essence [17, 18], modified gravity [19, 20, 21, 22, 23], the Dvali Gabadadze Porrati braneworld [24, 25], Chaplygin gas [26], holographic dark energy models [27, 28, 29], quantum cosmology [54], dynamical relaxation of the Cosmological Constant via membrane nucleation events of top forms [57] and multiple top forms [58]. For broad reviews see [30, 31].
Our goal here is to give a simple and intuitive origin for the cosmological term, and hence for dark energy, using only basic ideas that are compatible with but not restricted to string theory. The key input is the physics of extended objects with tension. We use this as a guide to show that a stress which is proportional to the metric arises naturally when a three dimensional medium with a constant volume tension fills space. We then argue that ordinary space itself can be modeled as such a medium. The resulting stress tensor has the same form as vacuum energy.
We begin with two parallel statements. For an infinitely extended fundamental three brane one finds,
$T_{\mu\nu(3)}=-T_{3}\,g_{\mu\nu}$ (1.10)
where $T_{3}$ is the brane tension with dimensions of energy density. If instead we treat space itself as an elastic three brane with an intrinsic volume tension $T_{s}$, then,
$T_{\mu\nu(s)}=-T_{s}\,g_{\mu\nu}$ (1.11)
Both expressions have the same structure as the vacuum stress tensor discussed earlier. The second statement is stronger conceptually because it identifies space itself as the source of the vacuum energy. We will analyze how a geometric tension of space and a four form sequestering structure interact with the vacuum sector and we will show that while matter vacuum energy can be neutralized by global constraints, a purely geometric NG-like tension cannot be removed. The residual cosmological term is then a geometric constant that we denote by $T_{s}$ and we identify it with Dark Energy.
2 Modeling Space as an Elastic Membrane
2.1 Review of Nambu–Goto Dynamics for Strings and Branes
A fundamental string in $1+1$ dimensions has worldsheet action,
$S_{1}=-T_{1}\int d^{2}\xi\,\sqrt{-\det(\gamma_{\alpha\beta})}$ (2.1)
with worldsheet coordinates $\xi^{\alpha}=(\tau,\sigma)$, string tension $T_{1}$ and induced metric,
$\gamma_{\alpha\beta}=\frac{\partial X^{\mu}}{\partial\xi^{\alpha}}\frac{\partial X^{\nu}}{\partial\xi^{\beta}}g_{\mu\nu}$ (2.2)
For a $d$ brane in (d+1)(d{+}1) dimensions the natural generalization is [32, 33, 34],
$S_{d}=-T_{d}\int d^{d+1}\xi\,\sqrt{(-1)^{d}\det(\gamma_{\alpha\beta})}$ (2.3)
where $T_{d}$ has dimensions of energy density in $d$ spatial dimensions. For $d\geq 3$ it is common to call $T_{d}$ a volume tension, since it generalizes surface tension to higher codimension.
We will specifically focus on the $d=3$ case,
$S_{3}=-T_{3}\int d^{4}\xi\,\sqrt{-\det(\gamma_{\alpha\beta})}$ (2.4)
2.2 Stress Tensor of an Infinitely Extended Three Brane
The stress tensor that couples to the ambient metric is defined by,
$T^{\mu\nu}_{(3)}(x)=\frac{2}{\sqrt{-g(x)}}\,\frac{\delta S_{3}}{\delta g_{\mu\nu}(x)}$ (2.5)
To vary $S_{3}$, note the standard determinant identity,
$\delta\sqrt{-\det\gamma}=\frac{1}{2}\sqrt{-\det\gamma}\,\gamma^{\alpha\beta}\delta\gamma_{\alpha\beta}$ (2.6)
Using $\delta\gamma_{\alpha\beta}=\partial_{\alpha}X^{\mu}\partial_{\beta}X^{\nu}\,\delta g_{\mu\nu}$, one finds,
$\frac{\delta S_{3}}{\delta g_{\mu\nu}(x)}=-\frac{T_{3}}{2}\int d^{4}\xi\,\sqrt{-\det\gamma}\,\gamma^{\alpha\beta}\,\frac{\partial X^{\mu}}{\partial\xi^{\alpha}}\frac{\partial X^{\nu}}{\partial\xi^{\beta}}\,\delta^{(4)}\big(x-X(\xi)\big)$ (2.7)
Thus,
$T^{\mu\nu}_{(3)}(x)=-\frac{T_{3}}{\sqrt{-g(x)}}\int d^{4}\xi\,\sqrt{-\det\gamma}\,\gamma^{\alpha\beta}\,\frac{\partial X^{\mu}}{\partial\xi^{\alpha}}\frac{\partial X^{\nu}}{\partial\xi^{\beta}}\,\delta^{(4)}\big(x-X(\xi)\big)$ (2.8)
For an infinitely extended brane the natural gauge choice is the static gauge. By choosing this gauge, we are basically identifying the time coordinate $\tau$ and the spatial coordinates $\sigma^{1},\sigma^{2},\sigma^{3}$ on the brane’s worldsheet with the coordinate time $t$ and spatial coordinates $x,y$ and $z$ of the target spacetime. In essence, the spacetime embedding coordinates of the brane simply become $X^{0}=\tau$ and $X^{i}=\sigma^{i}$ for $i=1,2,3$. Therefore, in the static gauge, we have,
$\frac{\partial X^{\mu}}{\partial\xi^{\alpha}}=\delta^{\mu}_{\alpha}$ (2.9)
where $\delta^{\mu}_{\alpha}$ is the Kronecker delta. Furthermore, equation (2.2) dictates that the induced metric $\gamma_{\alpha\beta}$ on the brane world-volume is related to the target spacetime metric $g_{\mu\nu}$ as,
$\gamma_{\alpha\beta}=g_{\mu\nu}\frac{\partial X^{\mu}}{\partial\xi^{\alpha}}\frac{\partial X^{\nu}}{\partial\xi^{\beta}}$ (2.10)
Substituting (2.9) in the equation above, we get,
$\gamma_{\alpha\beta}=g_{\mu\nu}\ \delta^{\mu}_{\alpha}\ \delta^{\nu}_{\beta}$ (2.11)
which simply reduces to,
$\gamma_{\alpha\beta}=g_{\alpha\beta}$ (2.12)
Therefore, we see that the induced metric $\gamma_{\alpha\beta}$ on the $3$-brane world-volume is equal to the target spacetime metric $g_{\alpha\beta}$. Furthermore, the equality above directly implies that $\sqrt{-\det(\gamma_{\alpha\beta})}=\sqrt{-g}$ and $\gamma^{\alpha\beta}=g^{\alpha\beta}$. This simplifies the expression for the $3$-brane stress energy tensor $T^{\mu\nu}_{(3)}$ (equation (2.8)) to,
$T^{\mu\nu}_{(3)}=-T_{3}\int d^{4}\xi\ g^{\alpha\beta}\frac{\partial X^{\mu}}{\partial\xi^{\alpha}}\frac{\partial X^{\nu}}{\partial\xi^{\beta}}\delta^{(3+1)}(x-X^{\mu}(\xi))$ (2.13)
In order to evaluate the expression above, we can re-write it as,
$T^{\mu\nu}_{(3)}=-T_{3}\int_{-\infty}^{\infty}d\tau\int_{-\infty}^{\infty}d^{3}\sigma\ g^{\alpha\beta}\frac{\partial X^{\mu}}{\partial\xi^{\alpha}}\frac{\partial X^{\nu}}{\partial\xi^{\beta}}\delta(x^{0}-\tau)\prod_{1}^{3}\delta(x^{i}-\sigma^{i})$ (2.14)
which allows us to straightforwardly perform the integral with respect to $\tau$ and obtain,
$T^{\mu\nu}_{(3)}=-T_{3}\int_{-\infty}^{\infty}d^{3}\sigma\ g^{\alpha\beta}\frac{\partial X^{\mu}}{\partial\xi^{\alpha}}\frac{\partial X^{\nu}}{\partial\xi^{\beta}}\prod_{1}^{3}\delta(x^{i}-\sigma^{i})$ (2.15)
Furthermore, given that we have assumed that the brane is infinitely large, we have,
$\int_{-\infty}^{\infty}d^{3}\sigma\prod_{1}^{3}\delta(x^{i}-\sigma^{i})=1$
since $x^{i}$ is always within the range of $\sigma^{i}$ $\forall\ i$. Therefore we get,
$T^{\mu\nu}_{(3)}=-T_{3}\ g^{\alpha\beta}\frac{\partial X^{\mu}}{\partial\xi^{\alpha}}\frac{\partial X^{\nu}}{\partial\xi^{\beta}}$ (2.16)
Using equation (2.9) in the equation above, we obtain,
$T^{\mu\nu}_{(3)}=-T_{3}\ g^{\alpha\beta}\ \delta^{\mu}_{\alpha}\ \delta^{\nu}_{\beta}$ (2.17)
which gives,
$T^{\mu\nu}_{(3)}=-T_{3}\ g^{\mu\nu}$ (2.18)
and finally, we obtain,
$T_{\mu\nu(3)}=-T_{3}\ g_{\mu\nu}$ (2.19)
as claimed in Sec. 1. This matches the expected units because $T_{3}$ is an energy density. Comparing with equation (1.5) in the Introduction we see that a uniform three brane tension $T_{3}$ behaves gravitationally like a vacuum energy density $\rho_{\mathrm{vac}}c^{2}$.
2.3 Space as the Three Brane
The previous derivation invites a sharper hypothesis. Instead of positing a separate embedded brane, we may treat space itself as a three dimensional elastic medium with an intrinsic volume tension $T_{s}$. The worldvolume of space is the physical spacetime and its worldvolume metric is simply $g_{\mu\nu}$. The Nambu–Goto like action for space is then,
$S_{s}=-T_{s}\int d^{4}x\,\sqrt{-g}$ (2.20)
This has the same dimensions as equation (2.4) with $T_{3}$ replaced by $T_{s}$. It also perfectly mirrors the cosmological piece of the Einstein–Hilbert action. The latter reads,
$S_{\mathrm{EH}}=\int d^{4}x\sqrt{-g}\left(\frac{c^{4}}{16\pi G}\,(R-2\Lambda)+\mathcal{L}_{M}\right)$ (2.21)
in which the cosmological piece is,
$S_{\Lambda}=-\frac{\Lambda c^{4}}{8\pi G}\int d^{4}x\sqrt{-g}$ (2.22)
Using $\Lambda=8\pi G\,\rho_{\mathrm{vac}}/c^{2}$ in the equation above, we get,
$S_{\Lambda}=-\rho_{\mathrm{vac}}c^{2}\int d^{4}x\sqrt{-g}$ (2.23)
Hence, dimensionally $\rho_{\mathrm{vac}}c^{2}]=[T_{s}]=[T_{3}]$, and $S_{s}$ given by equation (2.20) exactly mimics $S_{\Lambda}$ at the level of local dynamics.
We now look at the stress tensor that follows from the elastic action of space equation (2.20). On varying $S_{s}$, we obtain,
$\delta S_{s}=-T_{s}\int d^{4}x\,\delta\sqrt{-g}=-\frac{T_{s}}{2}\int d^{4}x\,\sqrt{-g}\,g^{\mu\nu}\delta g_{\mu\nu}$ (2.24)
Therefore,
$\frac{2}{\sqrt{-g}}\,\frac{\delta S_{s}}{\delta g_{\mu\nu}}=T_{s}\,g^{\mu\,\nu}$ (2.25)
But, the left hand side is the definition of the stress tensor for space, that is,
$T^{\mu\nu}_{(s)}=\frac{2}{\sqrt{-g}}\,\frac{\delta S_{s}}{\delta g_{\mu\nu}}$ (2.26)
Therefore, we have,
$T^{\mu\nu}_{(s)}=-T_{s}\,g^{\mu\nu}$ (2.27)
or that,
$T_{\mu\nu(s)}=-T_{s}\,g_{\mu\nu}$ (2.28)
Thus the elastic model of space reproduces the vacuum stress tensor with the identification $T_{s}=\rho_{\mathrm{vac}}c^{2}$.
Therefore, the Nambu–Goto structure ensures that a uniform volume tension produces a stress proportional to the metric. For an infinitely extended three brane this gives $T_{\mu\nu(3)}=T_{3}g_{\mu\nu}$. When we elevate space itself to the role of the three brane with intrinsic volume tension $T_{s}$, the same computation yields $T_{\mu\nu(s)}=-T_{s}g_{\mu\nu}$. Hence the cosmological term can be interpreted as the elastic volume tension of space. In the next section we will couple this picture to a four form sequestering framework and show that while matter vacuum energy can be neutralized by global constraints, a geometric volume tension cannot be removed. The residual dark energy is then precisely this $T_{s}$.
It is important to mention here that we do not assume that physical space is embedded in a higher dimensional ambient manifold. The elastic membrane language in this paper is an intrinsic description and all observables are built from the intrinsic metric and its derived tensors. Introducing an embedding is possible as a change of variables but it adds redundant structure that is not needed for the dynamics we study. An embedding also introduces an arbitrary ambient metric that does not enter any measurement performed by observers confined to the physical spacetime. Our equations of motion are fully determined by intrinsic geometry, so the theory remains sensible without any reference to extrinsic data. For these reasons we avoid committing to an embedding and keep the framework intrinsic from the outset.
3 A No-Geometric-Neutralization Theorem for Sequestering Four–Forms
We now work in natural units $\hbar=c=1$ in $3+1$ dimensions and introduce the sequester action which includes the usual Einstein–Hilbert term, a fundamental $3$-form Nambu-Goto action for space (essentially treating it as an elastic membrane), a matter sector that couples to a globally scaled metric through a single global variable $\lambda$ and two four-form sectors $F_{4}$ and $\hat{F}_{4}$ that enforce global constraints, see [61, 62, 63, 64, 65] for the detailed framework. The action has the following form,
$S=$ $\int d^{4}x\,\sqrt{-g}\,\Big(\frac{M_{*}^{2}}{2}\,R-T_{s}-\Lambda\Big)+\int d^{4}x\,\sqrt{-g}\,L_{m}(\lambda^{-2}\,g_{\mu\nu},\psi)$
$+\sigma\Big(\frac{\Lambda}{\mu^{4}}\Big)\,\int F_{4}+\hat{\sigma}\Big(\frac{\lambda^{4}}{\mu^{4}}\Big)\,\int\hat{F}_{4}$ (3.1)
with $M_{*}^{2}=1/8\pi G$. Note that we are working in a convention where the units of $\Lambda$ are $\Lambda]=\rm mass^{4}$ (energy density). The matter fields are collected into $\psi$ and the scale $\mu$ is a fixed reference scale. A note on notation, content of the action and minimal assumptions are in order,
a) We work with a Lorentzian metric $g_{\mu\nu}$ of signature (+++)(-+++). The Levi–Civita connection is metric compatible and curvature tensors have the usual definitions.
b) The total action includes the following pieces:
$\bullet$ Gravity: $\int\sqrt{-g}\,\frac{M_{*}^{2}}{2}\,R$.
$\bullet$ Geometric tension of space: $-\,T_{s}\int\sqrt{-g}$. One may think of $T_{s}$ as a constant that multiplies the physical spacetime volume. It is not a field and has no kinetic term and it only encodes an NG-like tension of the underlying elastic space. It is purely geometric in the sense that it couples only through $\sqrt{-g}$.
$\bullet$ A bare cosmological variable $\Lambda$ that appears as $-\Lambda\int\sqrt{-g}$.
$\bullet$ Matter with a global Weyl scaling through $\lambda$:
$S_{m}[g,\psi;\lambda]\equiv\int\sqrt{-g}\,L_{m}(\lambda^{-2}\,g_{\mu\nu},\psi)$ (3.2)
The factor $\lambda^{-2}g_{\mu\nu}$ ensures that the matter sector sees a metric that is globally Weyl related to $g_{\mu\nu}$. The variable $\lambda$ is not a field with spacetime dependence rather it is a single global variable. We will see this from the four-form sector later.
$\bullet$ Two top-form sectors,
$S_{4\text{-forms}}=\sigma\Big(\frac{\Lambda}{\mu^{4}}\Big)\int F_{4}+\hat{\sigma}\Big(\frac{\lambda^{4}}{\mu^{4}}\Big)\int\hat{F}_{4}$ (3.3)
where $F_{4}=dA_{3}$ and $\hat{F}_{4}=d\hat{A}_{3}$ with $A$ and $\hat{A}$ being three form potentials (see [54, 55, 57, 58] for a working understanding of three forms). The integrals $\int F_{4}$ and $\int\hat{F}_{4}$ do not use the volume factor $\sqrt{-g}$. In essence, they are purely topological and their values depend on global fluxes and boundary conditions, not on the local metric.
c) Assumptions about the functions $\sigma$ and $\hat{\sigma}$: We assume that they are smooth and monotone enough so that the equations we derive can be inverted to view $\Lambda$ and $\lambda$ as global constants fixed by fluxes. We do not need stronger properties at this stage. We also implicitly assume boundary conditions that render the variational problem well posed for $A_{3}$ and $\hat{A}_{3}$.
d) Variation conventions: Variation with respect to $g^{\mu\nu}$ defines the total stress tensor as usual. Variation with respect to $\psi$ gives matter equations of motion that we do not need explicitly. Variation with respect to $A_{3}$ and $\hat{A}_{3}$ gives four-form equations that enforce global constraints. Variation with respect to $\Lambda$ and $\lambda$ gives global relations because the four-forms enter as Lagrange multipliers as we will demonstrate below.
In all, this is the standard sequestering setup except there is an additional NG-like term for space.
We now proceed to show that under a constant shift of the matter Lagrangian $L_{m}$ by $\Delta V$ the combination $T^{(m)}_{\mu\nu}-\Lambda g_{\mu\nu}$ is invariant point-wise . The geometric tension $T_{s}$ is untouched by the four-form constraints because it does not couple to $\lambda$ and does not appear in the matter trace. Therefore the effective cosmological term left in the local Einstein equations equals $T_{s}$. A four-form sequestering mechanism of this type cannot neutralize the NG-like geometric tension of space.
Proof:
1.) Varying the action with respect to the three-form potentials:
Consider the term $\sigma(\Lambda/\mu^{4})\int F_{4}$, where $F_{4}=dA_{3}$. Under a variation of $A_{3}$ with compact support we have,
$\delta\left(\int\sigma\,F_{4}\right)=\int\sigma\,\delta F_{4}=\int\sigma\,d(\delta A_{3})=\int d\big(\sigma\,\delta A_{3}\big)-\int d\sigma\wedge\delta A_{3}$ (3.4)
The boundary term is $\int d(\sigma\,\delta A_{3})$. We set it to zero by appropriate boundary conditions. The remaining term is $\int d\sigma\wedge\delta A_{3}$ and given that $\delta A_{3}$ is arbitrary in the interior, the Euler–Lagrange equation is,
$d\sigma\Big(\frac{\Lambda}{\mu^{4}}\Big)=0$ (3.5)
Now because $\sigma$ depends on spacetime only through $\Lambda$, this implies that $d(\Lambda/\mu^{4})=0$ provided $\sigma^{\prime}\neq 0$. Therefore $\Lambda$ is constant in spacetime.
Repeating the same steps for $\hat{\sigma}(\lambda^{4}/\mu^{4})\int\hat{F}_{4}$, we have after varying with respect to $\hat{A}_{3}$,
$d\hat{\sigma}\Big(\frac{\lambda^{4}}{\mu^{4}}\Big)=0$ (3.6)
Assuming $\hat{\sigma}^{\prime}\neq 0$ we conclude that $d(\lambda^{4}/\mu^{4})=0$. Hence $\lambda$ is constant in spacetime and since $\lambda>0$ we can take $\lambda$ itself to be a global constant.
Therefore, the four-form sectors force $\Lambda$ and $\lambda$ to be global constants. That is, they do not propagate and they are numbers fixed by boundary data and fluxes.
2.) Varying the action with respect to the metric and identifying the local Einstein equations:
We now compute the variation of each metric dependent term with respect to $g^{\mu\nu}$ where we will be using the identity $\delta\sqrt{-g}=-\dfrac{1}{2}\sqrt{-g}\,g_{\mu\nu}\,\delta g^{\mu\nu}$. The variation of the Einstein–Hilbert part gives the usual $M_{*}^{2}G_{\mu\nu}$ on the left hand side of the field equations.
The variation of $T_{s}\int\sqrt{-g}$ gives,
$\delta\big(-T_{s}\int\sqrt{-g}\big)=-T_{s}\int\delta\sqrt{-g}=\frac{T_{s}}{2}\int\sqrt{-g}\,g_{\mu\nu}\,\delta g^{\mu\nu}$ (3.7)
This means the geometric tension contributes a stress tensor equal to $T_{s}g_{\mu\nu}$ with a minus sign when moved to the right hand side of Einstein equations.
The variation of $\Lambda\int\sqrt{-g}$ is identical in structure and it gives a negative contribution $\Lambda g_{\mu\nu}$ to the right hand side of the field equations as well.
Now we vary the matter action with respect to the metric, which reads,
$S_{m}[g,\psi;\lambda]=\int\sqrt{-g}\,\,L_{m}(\lambda^{-2}\,g,\psi)$ (3.8)
We define the physical matter stress tensor by the standard definition,
$T^{(m)}_{\mu\nu}\equiv-\frac{2}{\sqrt{-g}}\,\frac{\delta S_{m}}{\delta g^{\mu\nu}}$ (3.9)
The above definition ensures that the matter equations are consistent with diffeomorphism invariance and with the Bianchi identities. The full variation of the matter part yields,
$\delta S_{m}=-\frac{1}{2}\int\sqrt{-g}\,T^{(m)}_{\mu\nu}\,\delta g^{\mu\nu}$ (3.10)
Collecting all terms, the total variation with respect to $g^{\mu\nu}$ gives,
$\delta S=\int\sqrt{-g}\,\Big[\frac{M_{*}^{2}}{2}\,G_{\mu\nu}+\frac{1}{2}\big(T_{s}+\Lambda\big)g_{\mu\nu}-\frac{1}{2}\,T^{(m)}_{\mu\nu}\Big]\,\delta g^{\mu\nu}$ (3.11)
Stationarity for arbitrary $\delta g^{\mu\nu}$ gives,
$\frac{M_{*}^{2}}{2}\,G_{\mu\nu}+\frac{1}{2}\big(T_{s}+\Lambda\big)g_{\mu\nu}-\frac{1}{2}\,T^{(m)}_{\mu\nu}=0$ (3.12)
Multiplying by 2, the local Einstein field equations read,
$M_{*}^{2}\,G_{\mu\nu}=T^{(m)}_{\mu\nu}-\big(T_{s}+\Lambda\big)\,g_{\mu\nu}$ (E)
It is important to note that the four-form terms do not contribute to the metric equations. They are metric independent in the form we wrote, since $\int F_{4}$ and $\int\hat{F}_{4}$ do not carry $\sqrt{-g}$ or $g^{\mu\nu}$. Hence they add no local stress tensor.
So, we now have the local field equations where we see that the geometric tension $T_{s}$ and the variable $\Lambda$ both appear as volume terms and therefore enter as $(\Lambda+T_{s})\,g_{\mu\nu}$. The matter sector contributes its usual stress tensor $T^{(m)}_{\mu\nu}$.
3.) Varying the action with respect to $\Lambda$:
Only two parts of the action (3.1) depend on $\Lambda$. The volume term $\Lambda\int\sqrt{-g}$ and the top-form term $\sigma(\Lambda/\mu^{4})\int F_{4}$. Therefore, computing the variation and treating $\Lambda$ as a global number, we find,
$\delta S\big|_{\Lambda}=-\delta\Lambda\int d^{4}x\,\sqrt{-g}+\sigma^{\prime}\Big(\frac{\Lambda}{\mu^{4}}\Big)\,\frac{\delta\Lambda}{\mu^{4}}\,\int F_{4}$ (3.13)
Setting the equation above this to zero for arbitrary $\delta\Lambda$, we obtain the first global constraint,
$\int d^{4}x\,\sqrt{-g}=\frac{\sigma^{\prime}\left(\Lambda/\mu^{4}\right)}{\mu^{4}}\,\int F_{4}$ (C1)
This equation tells us that the spacetime four-volume is proportional to the four-form flux weighted by $\sigma^{\prime}$. Note that it does not fix the local metric rather it only relates global numbers.
With this, we have one global relation that involves $\Lambda$ but no local fields directly.
4.) Varying the action with respect to $\lambda$:
Again, only two parts of the action (3.1) depend on $\lambda$, which are the matter action $S_{m}$ and the top-form term $\hat{\sigma}(\lambda^{4}/\mu^{4})\int\hat{F}_{4}$. First consider the matter action which we compute $\delta S_{m}$ under a small variation of the global $\lambda$.
$\delta S_{m}\bigg|_{\lambda}=\int\sqrt{-g}\,\frac{\partial L_{m}(\lambda^{-2}\,g,\,\psi)}{\partial(\lambda^{-2}g_{\mu\nu})}\,\delta(\lambda^{-2}g_{\mu\nu})$ (3.14)
Let $h_{\mu\nu}\equiv\lambda^{-2}g_{\mu\nu}$ and its variation under a variation of $\lambda$ at fixed $g$ is,
$\delta h_{\mu\nu}=\delta(\lambda^{-2})\,g_{\mu\nu}=-2\,\lambda^{-3}\,\delta\lambda\,g_{\mu\nu}$ (3.15)
So we have,
$\int\sqrt{-g}\;\frac{\partial L_{m}(h_{\mu\nu},\,\psi)}{\partial(h_{\mu\nu})}\;\delta(h_{\mu\nu})=-2\,\int\sqrt{-g}\;\lambda^{-3}\,\frac{\partial L_{m}(h_{\mu\nu},\,\psi)}{\partial(h_{\mu\nu})}\,g_{\mu\nu}\,\delta\lambda$ (3.16)
In order to obtain $\dfrac{\partial L_{m}(h_{\mu\nu},\,\psi)}{\partial(h_{\mu\nu})}\,g_{\mu\nu}$ , we first vary $S_{m}$ with respect to the metric tensor $g_{\mu\nu}$, which gives,
$\delta S_{m}=-\int\sqrt{-g}\;\lambda^{-2}\,\frac{\partial L_{m}(h_{\mu\nu},\psi)}{\partial(h_{\mu\nu})}g_{\alpha\mu}\,g_{\beta\nu}\delta g^{\mu\nu}$ (3.17)
and use the definition of the matter stress energy tensor (3.9), which gives,
$T_{\mu\nu}^{(m)}=2\,\lambda^{-2}\,\frac{\partial L_{m}(h_{\mu\nu},\psi)}{\partial(h_{\mu\nu})}g_{\alpha\mu}\,g_{\beta\nu}$ (3.18)
Taking the trace, we find,
$T^{(m)}=g^{\mu\nu}\,T_{\mu\nu}^{(m)}=2\,\lambda^{-2}\,\frac{\partial L_{m}(h_{\mu\nu},\psi)}{\partial(h_{\mu\nu})}g_{\mu\nu}$ (3.19)
Solving for $\dfrac{\partial L_{m}(h_{\mu\nu},\,\psi)}{\partial(h_{\mu\nu})}\,g_{\mu\nu}$, we obtain,
$\dfrac{\partial L_{m}(h_{\mu\nu},\,\psi)}{\partial(h_{\mu\nu})}g_{\mu\nu}=\frac{\lambda^{2}\,T^{(m)}}{2}$ (3.20)
Substituting into equation (3.16), we get,
$\int\sqrt{-g}\;\frac{\partial L_{m}(h_{\mu\nu},\,\psi)}{\partial(h_{\mu\nu})}\;\delta(h_{\mu\nu})=-\frac{\delta\lambda}{\lambda}\int\sqrt{-g}\,T^{(m)}$ (3.21)
Using (3.21) in (3.14), we obtain,
$\delta S_{m}\bigg|_{\lambda}=-\frac{\delta\lambda}{\lambda}\int\sqrt{-g}\,T^{(m)}$ (3.22)
Now we vary the top-form term $\hat{\sigma}(\lambda^{4}/\mu^{4})\int\hat{F}_{4}$, which gives,
$\delta\Big[\hat{\sigma}\!\Big(\frac{\lambda^{4}}{\mu^{4}}\Big)\int\hat{F}_{4}\Big]=\hat{\sigma}^{\prime}\Big(\frac{\lambda^{4}}{\mu^{4}}\Big)\,\frac{4\lambda^{3}\,\delta\lambda}{\mu^{4}}\,\int\hat{F}_{4}$ (3.23)
Writing both pieces together, that is equations (3.21) and (3.23), the full variation of $S_{m}$ with respect to the global $\lambda$ then reads,
$\frac{\delta\lambda}{\lambda}\,\int\sqrt{-g}\,T^{(m)}+\hat{\sigma}^{\prime}\Big(\frac{\lambda^{4}}{\mu^{4}}\Big)\,\frac{4\lambda^{3}\,\delta\lambda}{\mu^{4}}\,\int\hat{F}_{4}=0$ (3.24)
Canceling $\delta\lambda$ which is arbitrary but not zero and multiplying by $\lambda$, we obtain the second global constraint,
$\int d^{4}x\,\sqrt{-g}\,T^{(m)}=\frac{4\,\lambda^{4}}{\mu^{4}}\,\hat{\sigma}^{\prime}\left(\frac{\lambda^{4}}{\mu^{4}}\right)\,\int\hat{F}_{4}$ (C2)
We now have a second global relation that connects the spacetime integral of the matter trace to the flux of $\hat{F}_{4}$. The geometric tension $T_{s}$ does not appear because it does not depend on $\lambda$.
5.) Eliminating fluxes to relate $\Lambda$ to the spacetime average of $T^{(m)}$:
Equations (C1) and (C2) contain the fluxes $\int F_{4}$ and $\int\hat{F}_{4}$. These fluxes are constant numbers fixed by boundary conditions. We define the spacetime average of a quantity $X$ as
$\langle X\rangle=\dfrac{\int\sqrt{-g}\,X}{\int\sqrt{-g}}$ (3.25)
Combining (C1) and (C2) we may then write,
$\langle T^{(m)}\rangle=4\,\lambda^{4}\,\frac{\hat{\sigma}^{\prime}(\lambda^{4}/\mu^{4})}{\sigma^{\prime}(\Lambda/\mu^{4})}\,\frac{\int\hat{F}_{4}}{\int F_{4}}$ (3.26)
The right hand side is a pure number fixed by the values of the global variables and by the flux ratio. In the canonical sequestering normalization this relation is equivalent to,
$\Lambda=\frac{1}{4}\,\langle T^{(m)}\rangle$ (3.27)
This is the standard global identity in four-form sequestering and follows from the homogeneity in $\lambda$ together with the two Lagrange multiplier equations. Therefore, we have a global identity that expresses $\Lambda$ as one quarter of the spacetime average of the matter trace.
6.) Performing the constant shift test and identify the invariant local combination:
Consider a rigid shift of the matter Lagrangian by a constant $\Delta V$. Equivalently, we may replace $L_{m}\to L_{m}-\Delta V$ so that the added term in the action is $\int\sqrt{-g}\,\lambda^{4}\Delta V$. Using equation (3.9), we see that this simply changes the matter stress $T^{(m)}_{\mu\nu}$ and trace of matter stress $T^{(m)}$ by,
$\Delta T^{(m)}_{\mu\nu}=-\Delta V\,g_{\mu\nu},\qquad\Delta T^{(m)}=-4\,\Delta V$ (3.28)
The global relation then yields,
$\Delta\Lambda=\frac{1}{4}\,\Delta\langle T^{(m)}\rangle=-\,\Delta V$ (3.29)
Therefore the matter-vacuum local combination that appears in the Einstein equations is invariant,
$T^{(m)}_{\mu\nu}-\Lambda g_{\mu\nu}\;\to\;\big(T^{(m)}_{\mu\nu}-\Delta Vg_{\mu\nu}\big)-\big(\Lambda-\Delta V\big)g_{\mu\nu}=T^{(m)}_{\mu\nu}-\Lambda g_{\mu\nu}$ (3.30)
All constant contributions from the matter sector cancel locally between $T^{(m)}_{\mu\nu}$ and $\Lambda g_{\mu\nu}$. This includes radiative corrections that enter as shifts of the vacuum energy in $L_{m}$.
7.) Identifying the leftover geometric NG-term:
Returning to the local Einstein equations (E),
$M_{*}^{2}\,G_{\mu\nu}=T^{(m)}_{\mu\nu}-\big(T_{s}+\Lambda\big)\,g_{\mu\nu}$ (3.31)
We rewrite the right hand side as,
$\big[T^{(m)}_{\mu\nu}-\Lambda g_{\mu\nu}\big]-T_{s}\,g_{\mu\nu}$ (3.32)
The bracket is invariant under constant shifts of the matter Lagrangian. Therefore any constant vacuum energy from the matter sector is neutralized by the four-form enforced value of $\Lambda$. The term $T_{s}g_{\mu\nu}$ is untouched by this mechanism because $T_{s}$ does not depend on $\lambda$ and does not appear in the global trace equation. It is a purely geometric constant and it contributes a cosmological term to the Einstein equations that cannot be removed or adjusted by the four-form constraints. Therefore the effective cosmological term that remains in the local field equations is $T_{s}$.
Thus, the neutralization does not act on $T_{s}$. The residual cosmological constant is $T_{s}$ as claimed.
8.) General Remarks:
A. Bianchi identities and conservation: Taking the divergence of both sides of the Einstein equations, the left hand side obeys $\nabla^{\mu}G_{\mu\nu}=0$ by the contracted Bianchi identity. The right hand side has $\nabla^{\mu}T^{(m)}_{\mu\nu}=0$ because the matter action is diffeomorphism invariant and couples minimally to $g_{\mu\nu}$. The term $(T_{s}+\Lambda)\,g_{\mu\nu}$ has zero divergence because $T_{s}$ and $\Lambda$ are global constants. Hence the equations are consistent with the Bianchi identities.
B. Role of the top forms in the stress tensor: In the action (3.1) the four-form integrals carry no metric dependence and therefore do not contribute any local stress. This is why the local equations contain only $T^{(m)}_{\mu\nu}$ and the volume terms. If one used a different four-form model that included metric dependent contractions of $F_{\mu\nu\rho\sigma}$ with itself, then the four-form would carry local stress and would behave like a dynamical cosmological constant. That is the Brown–Teitelboim mechanism. Our conclusion is specific to the sequestering setup where the four-forms serve as global Lagrange multipliers rather than as local energy carriers.
C. Radiative stability: Matter loop corrections shift the matter vacuum energy by constants of arbitrarily large size. These shifts are neutralized locally because the bracket $T^{(m)}_{\mu\nu}-\Lambda g_{\mu\nu}$ is invariant. Gravitational loops can renormalize the geometric constant $T_{s}$ since it multiplies the volume element. In an effective field theory description one treats $T_{s}$ as a renormalized parameter that must be fixed by observation. So the key point is that matter loops do not destabilize the observed value once $T_{s}$ is chosen. We will discuss this point at length in the next Section.
D. Scope of the theorem. The argument required only that the matter sector couples through a single global Weyl factor $\lambda$. It required that the geometric tension $T_{s}$ does not carry any explicit $\lambda$ dependence. If one added a non minimal coupling that made $T_{s}$ scale with $\lambda$, then the cancellation pattern would change and the result would not follow. Our setup excludes such couplings by assumption.
9.) Einstein Equations for a pure matter vacuum:
Consider a pure matter vacuum with $T^{(m)}_{\mu\nu}=-Vg_{\mu\nu}$ for some constant $V$. The trace is $T^{(m)}=-4\,V$ and the spacetime average of the trace equals $\langle T^{(m)}\rangle=-4V$. The global constraint gives $\Lambda=\dfrac{1}{4}\langle T^{(m)}\rangle=-V$. The matter-vacuum bracket therefore vanishes,
$T^{(m)}_{\mu\nu}-\Lambda g_{\mu\nu}]=0$ (3.33)
and the Einstein equations reduce to,
$M_{*}^{2}G_{\mu\nu}=-T_{s}\,g_{\mu\nu}$ (3.34)
The geometric term $T_{s}$ in turn receives correction from graviton self loops and graviton-matter loops as we show in the subsequent Section. Therefore, it can be naturally identified with the residual observed Dark Energy. This conclusion is consistent with the general statement of the theorem and also makes the no geometric sequestering mechanism very transparent.
$\square$
4 Gravitational EFT, Graviton Loops and the Residual Term $T_{s}$
Our goal is to present the low energy gravitational effective action in a way that keeps the discussion compact while preserving all structural information, see [36, 37]. We use the background field method [39, 40], expand only the metric around a fixed background and keep matter fields unexpanded in this section. Background diffeomorphism invariance then forces the one particle irreducible functional $\Gamma$ to be a sum of background covariant operators. This locks in the tensor structure of counterterms and of the finite Wilsonian terms. Sequestering relation will then isolate the surviving constant source as the geometric volume tension $T_{s}$.
We introduce a background metric $\bar{g}_{\mu\nu}$ and the fluctuation
$g_{\mu\nu}\equiv\bar{g}_{\mu\nu}+\kappa\,h_{\mu\nu},\qquad\kappa^{2}\equiv 32\pi G\equiv\frac{4}{M_{*}^{2}}$ (4.1)
The action is still given by (3.1),
$S[g,\psi;\Lambda,\lambda]=\int d^{4}x\,\sqrt{-g}\,\Big(\tfrac{M_{*}^{2}}{2}\,R-T_{s}-\Lambda\Big)+S_{m}[\lambda^{-2}g,\psi]+S_{\text{4-forms}}[\Lambda,\lambda]$ (4.2)
where the four form sector enforces the global constant constrains on $\Lambda$ and $\lambda$.
We couple a source $J^{\mu\nu}$ to the metric fluctuation and define the gauge fixed partition function,
$Z[\bar{g},J]=\int\mathcal{D}h\,\mathcal{D}c\,\mathcal{D}\bar{c}\,\mathcal{D}\psi\;e^{i\Big(S[\bar{g}+\kappa h,\psi]+S_{\text{gf}}[\bar{g},h]+S_{\text{gh}}[\bar{g},h,c,\bar{c}]+\int d^{4}x\,\sqrt{-\bar{g}}\,J^{\mu\nu}h_{\mu\nu}\Big)}$ (4.3)
where $S_{\text{gf}}[\bar{g},h]$ and $S_{\text{gh}}[\bar{g},h,c,\bar{c}]$ are the gravitational gauge fixing and ghost Lagrangians respectively as discussed below.
We now define the connected generator,
$W[\bar{g},J]\equiv-\,i\ln Z[\bar{g},J]$ (4.4)
The mean fluctuation field is therefore,
$\hat{h}_{\mu\nu}(x)\equiv\frac{1}{\sqrt{-\bar{g}}}\frac{\delta W[\bar{g},J]}{\delta J^{\mu\nu}(x)}$ (4.5)
We have introduced $\hat{h}$ in order to Legendre transform to the one-particle-irreducible functional. This produces vertex functions, clean Ward identities and the tadpole condition that fixes the background. The Legendre transform gives,
$\Gamma[\bar{g},\hat{h}]\equiv W[\bar{g},J]-\int d^{4}x\,\sqrt{-\bar{g}}\,J^{\mu\nu}\hat{h}_{\mu\nu},\qquad\frac{1}{\sqrt{-\bar{g}}}\frac{\delta\Gamma}{\delta\hat{h}_{\mu\nu}}=-J^{\mu\nu}$ (4.6)
Setting $J=0$ enforces the stationarity condition
$\left.\frac{\delta\Gamma[\bar{g},\hat{h}]}{\delta\hat{h}_{\mu\nu}}\right|_{J=0}=0$ (4.7)
By construction we choose the background so that $\hat{h}=0$ solves (4.7). The vacuum functional is then,
$\Gamma[\bar{g}]\equiv\Gamma[\bar{g},\hat{h}]\big|_{\hat{h}=0}$ (4.8)
This object generates the quantum-corrected field equations for the background metric.
We now fix diffeomorphisms with a background covariant gauge. A standard choice is de Donder gauge built from $\bar{\nabla}$, where $\bar{\nabla}$ (the covariant derivative defined on background metric $\bar{g}$) is the unique torsion–free, metric–compatible connection satisfying,
$\bar{\nabla}_{\rho}\bar{g}_{\mu\nu}=0,\qquad\bar{T}^{\rho}{}_{\mu\nu}\equiv\bar{\Gamma}^{\rho}{}_{\mu\nu}-\bar{\Gamma}^{\rho}{}_{\nu\mu}=0$ (4.9)
and where,
$\bar{\Gamma}^{\rho}{}_{\mu\nu}=\tfrac{1}{2}\,\bar{g}^{\rho\sigma}\big(\partial_{\mu}\bar{g}_{\nu\sigma}+\partial_{\nu}\bar{g}_{\mu\sigma}-\partial_{\sigma}\bar{g}_{\mu\nu}\big)$ (4.10)
Using de Donder’s gauge, we have,
$F_{\mu}^{\text{grav}}=\bar{\nabla}^{\nu}h_{\mu\nu}-\frac{1}{2}\bar{\nabla}_{\mu}h,\qquad\mathcal{L}_{\text{gf}}^{\text{grav}}=-\frac{1}{2\alpha_{g}}\,\sqrt{-\bar{g}}\,\bar{g}^{\mu\nu}F_{\mu}F_{\nu}$ (4.11)
The Faddeev–Popov ghost action is then [41],
$\Delta_{\ \nu}^{\mu}=\bar{\nabla}^{2}\delta^{\mu}_{\ \nu}+\bar{R}^{\mu}_{\ \nu},\qquad\mathcal{L}_{\text{gh}}^{\text{grav}}=-\sqrt{-\bar{g}}\,\bar{c}_{\mu}\,\Delta_{\ \nu}^{\mu}\,c^{\nu}$ (4.12)
Let $\xi^{\mu}(x)$ generate an infinitesimal diffeomorphism. Background diffeomorphisms act as ordinary diffeomorphisms on $\bar{g}$ and as tensorial Lie derivatives on all quantum fields and on the source $J$,
$\delta_{\text{bg}}\bar{g}_{\mu\nu}=\bar{\nabla}_{\mu}\xi_{\nu}+\bar{\nabla}_{\nu}\xi_{\mu},\qquad\delta_{\text{bg}}h_{\mu\nu}=\mathcal{L}_{\xi}h_{\mu\nu},$
$\delta_{\text{bg}}c^{\mu}=\mathcal{L}_{\xi}c^{\mu},\qquad\delta_{\text{bg}}\bar{c}_{\mu}=\mathcal{L}_{\xi}\bar{c}_{\mu},\qquad\delta_{\text{bg}}J^{\mu\nu}=\mathcal{L}_{\xi}J^{\mu\nu}$ (4.13)
Because $S_{\text{gf}}$ and $S_{\text{gh}}$ are built from $\bar{g}$ and $\bar{\nabla}$, they are invariant under $\delta_{\text{bg}}$. The classical action $S[\bar{g}+\kappa h,\psi]$ is diffeomorphism invariant by construction. Furthermore, we adopt a background-covariant regularization and measure, for instance dimensional regularization [38], so that $Z$ and $W$ are invariant when $J$ transforms as a background tensor. The Legendre transform then gives the Ward identity,
$\delta_{\text{bg}}\Gamma[\bar{g},\hat{h}]=0,\qquad\delta_{\text{bg}}\hat{h}_{\mu\nu}=\mathcal{L}_{\xi}\hat{h}_{\mu\nu}$ (4.14)
If we set $\hat{h}=0$, we obtain $\delta\,\Gamma[\bar{g}]=0$, which implies that $\Gamma[\bar{g}]$ is invariant under diffeomorphisms of $\bar{g}$ and therefore it can only be an integral of scalar densities built from $\bar{g}_{\mu\nu}$, the Riemann tensor of $\bar{g}$ and background covariant derivatives. This is the structural reason why the effective action organizes into background-covariant operators.
Equation (4.14) is the reason why the effective action admits an expansion in background covariant operators. Every local counterterm and every local Wilsonian term (see [51, 42, 43, 44]) is an integral of $\sqrt{-\bar{g}}$ times a scalar built from $\bar{R}$, $\bar{R}_{\mu\nu}$, $\bar{R}_{\mu\nu\rho\sigma}$, and $\bar{\nabla}$ and it takes the universal form for the local part,
$\Gamma^{\text{local}}[\bar{g}]=\int d^{4}x\,\sqrt{-\bar{g}}\,\Big[c_{0}(\mu)+\frac{1}{2}\,c_{2}(\mu)\,R+c_{4}^{(1)}(\mu)\,R^{2}+c_{4}^{(2)}(\mu)\,R_{\mu\nu}\,R^{\mu\nu}+\cdots\Big]$ (4.15)
The non-local part contains structures such as $R\log(-\bar{\Box}/\mu^{2})R$ [49, 45, 50]. These modify long distance kernels but do not act as constant background sources such as dark energy because they contain derivatives.
In particular, the coefficient $c_{0}(\mu)$ renormalizes the volume term, that is $T_{s}$, the coefficient $c_{2}(\mu)$ renormalizes the Einstein–Hilbert coupling (Newton’s constant $G$), the $c_{4}^{(i)}(\mu)$ term renormalizes curvature squared operators and higher terms contain more derivatives and are suppressed at low curvature.
The parameters $c_{n}(\mu)$ run with the renormalization scale $\mu$. We write,
$\mu\,\frac{d}{d\mu}\,c_{0}(\mu)=\beta_{0}(\mu),\qquad\mu\,\frac{d}{d\mu}\,c_{2}(\mu)=\beta_{2}(\mu),\qquad\mu\,\frac{d}{d\mu}\,c^{(i)}_{4}(\mu)=\beta^{(i)}_{4}(\mu),\quad\text{etc.}$ (4.16)
The precise values depend on the field content and on the gauge fixing and on the regulator.
We now combine tree level and loop corrected coefficients that multiply the same operators. We define renormalized parameters at the scale $\mu$ by
$T_{s}(\mu)$ $\equiv T_{s}^{\text{bare}}-c_{0}(\mu),$
$M_{*}^{2}(\mu)$ $\equiv M_{*\,\text{bare}}^{2}+c_{2}(\mu),$
$\alpha_{1,2}(\mu)$ $\equiv c_{4}^{(1,2)}(\mu)$ (4.17)
The local gravitational part of the effective action can therefore be written as,
$\Gamma^{\text{grav}}_{\text{local}}=\int d^{4}x\,\sqrt{-g}\,\Big(-T_{s}(\mu)+\frac{M_{*}^{2}(\mu)}{2}R+\alpha_{1}(\mu)R^{2}+\alpha_{2}(\mu)R_{\mu\nu}R^{\mu\nu}+\cdots\Big)$ (4.18)
Varying and adding the matter part gives the local field equations,
$M_{*}^{2}(\mu)\,G_{\mu\nu}=[T^{(m)}_{\mu\nu}-\Lambda\,g_{\mu\nu}]-T_{s}(\mu)\,g_{\mu\nu}+\mathcal{H}_{\mu\nu}[R^{2},\dots]$ (4.19)
where $\mathcal{H}_{\mu\nu}$ collects higher derivative tensors from curvature squared terms. These tensors are suppressed at low curvature and do not act like a constant source. The combination $T^{(m)}_{\mu\nu}-\Lambda g_{\mu\nu}$ is invariant under rigid shifts $L_{m}\to L_{m}-\Delta V$ in the sequestering setup, hence constant vacuum energy inside $S_{m}$ cancels out of the local equations.
So in the field equation (4.19), the net constant vacuum source after sequestering is
$\rho_{\rm vac}(\mu)\,g_{\mu\nu}=-T_{s}(\mu)\,g_{\mu\nu}$ (4.20)
At late times one can match this constant source to the observed dark energy density,
$\rho_{\Lambda}^{\rm obs}=T_{s}(\mu_{\rm IR})$ (4.21)
where $\mu_{\rm IR}$ is an infrared matching scale appropriate for cosmology. This identification fixes one linear combination of low energy parameters using data.
5 Discussion
We developed a minimal framework in which space is modeled as an elastic three membrane that fills spacetime. The dynamics are captured by an intrinsic Nambu–Goto volume term $S_{s}=-T_{s}\int\sqrt{-g}\,d^{4}x$. This fixes the geometric stress tensor to $T^{(s)}_{\mu\nu}=-T_{s}\,g_{\mu\nu}$ and identifies the membrane tension with a uniform vacuum energy of space. We then embedded this term into a matter vacuum sequester framework that employs global four form constraints to cancel constant contributions from the matter sector. The variation shows that the global variable adjusts to neutralize the matter vacuum energy while a purely geometric unit operator (the operator which behaves like vacuum energy) carries no dependence on that variable and is not in the cancellation channel. As a result the local field equations retain a residual contribution $T_{s}\,g_{\mu\nu}$. We interpret this as the effective cosmological term.
Subsequently we performed the path integral over the gravitational fluctuation $h_{\mu\nu}$ and over the associated Faddeev–Popov ghosts only. We did not integrate over matter fluctuations so as to not “recompute” the pure matter loops. Their effect is assumed to have already been absorbed into the renormalized matter functional $S_{m}[\lambda^{-2}g,\psi]$ at the matching scale $\mu$.
The outcome was a background diffeomorphism invariant one particle irreducible action $\Gamma[\bar{g}]$ with the structure shown in equation (4.15). The local part is an infinite covariant tower built from the curvature and its derivatives. The non-local part contains form factors that never reduce to a pure volume term and so do not influence vacuum energy. The geometric volume tension $T_{s}(\mu)$ arises from the renormalized coefficient of the unit operator and enters the field equations. This separation makes clear which ingredients can shift $T_{s}(\mu)$ and which ingredients only change higher derivative responses of the geometry.
If one wishes to include matter-graviton mixed loops explicitly, one can extend the path integral to the matter fields fluctuations as well. The background field method and general covariance then enforce the same covariant operator basis that appears in equation (4.15). Mixed matter-graviton diagrams that contain at least one internal graviton line are local in the ultraviolet. They therefore renormalize the same coefficients that multiply curvature invariants as in equation (4.15). No new local scalar of lower canonical mass dimension exists beyond the unit operator and the Ricci scalar term. As a result the only change is in the numerical values of the Wilson coefficients.
Concretely, mixed matter–graviton loops shift the same $c_{0}(\mu)$ and therefore in turn shift $T_{s}(\mu)\equiv T_{s}^{\rm bare}-c_{0}(\mu)$. They also shift $c_{2}(\mu)$ which renormalizes the constant $G$ along with shifting the coefficients of curvature squared invariants and beyond. The tensorial form of each contribution is fixed by background diffeomorphisms and does not depend on the details of the gauge fixing. It can further be arranged so that pure matter functional $S_{m}[\lambda^{-2}g,\psi]$ will still continue to encode matter self interactions and pure matter loops at the chosen matching scale. The combined result is a single covariant effective action with renormalized coefficients $c_{0}(\mu),c_{1}(\mu),\cdots$ that capture both gravitational and mixed contributions in a uniform way. Therefore, the generalization to mixed matter-graviton graphs is straightfoward.
With this in mind, one can then conveniently regard $T_{s}(\mu)$ as the fully renormalized geometric tension that already includes contributions from pure graviton self loops and from mixed matter–graviton loops because as mentioned before, extending the path integral from metric fluctuations to metric plus matter fluctuations only shifts the Wilson coefficients in the same covariant operator basis. The definition of $T_{s}(\mu)$ as the renormalized coefficient of the unit operator therefore captures all such shifts in a single parameter.
One might try to redefine $\Lambda^{\prime}\equiv\Lambda-T_{s}$ so that the sequestering constraint cancels both the matter vacuum energy and the geometric volume term. This identification is physically not possible because in the standard sequestering framework which we are considering the global variables and four-forms are wired so that the variational constraint removes constant shifts of the matter Lagrangian, while the purely gravitational sector is inert under that constraint. An independent geometric unit operator $T_{s}\int\sqrt{-g}$ is therefore not neutralized unless one explicitly extends the mechanism to couple the same global sector to that operator. To also cancel $T_{s}$, one would need a different theory, one which couples a new rigid variable to a geometric density. This represents a different sequestering scheme and whereas we are working in the minimal framework which requires just two global Lagrange multipliers $\lambda$ and $\Lambda$. One can also see why such an identification is not possible in the minimal case because the operator $T_{s}\int\sqrt{-g}$ is blind to $\lambda$ and therefore cannot contribute to the $\lambda$ constrain.
So it is correct to keep $\Lambda$ as the variable that enforces the global matter constraint and we keep $T_{s}(\mu)$ as an intrinsic geometric coupling that lives in the gravitational effective action. In the minimal sequester framework, the added advantage of having an un-sequestered geometric term is that after fixing covariant operator basis using the background field method, the term $T_{s}(\mu)$ automatically emerges as the renormalized coefficient of the unit operator after graviton and mixed matter–graviton loops are included. Nonlocal form factors appear but they do not produce a pure volume term and therefore do not modify this identification. The constant source in the field equations can therefore naturally be identified as $\rho_{\rm cav}(\mu)=T_{s}(\mu)$. At an infrared scale $\mu_{\rm IR}$ appropriate for cosmology one can then match $\rho_{\rm vac}(\mu_{\rm IR})$ to the observed dark energy density. This retains the virtue of sequestering for matter vacuum energy while leaving a natural geometric volume term that can account for the residual dark energy. It also avoids promoting geometric couplings to global multipliers and keeps conservation and locality manifest through the Bianchi identity.
6 Conclusion
We have modeled space as an elastic three membrane and identified its uniform volume tension with a geometric unit operator that contributes a stress proportional to the metric. The Nambu–Goto term $S_{s}=-T_{s}\int\sqrt{-g}\,d^{4}x$ fixes $T^{(s)}_{\mu\nu}=-T_{s}g_{\mu\nu}$ and allows a direct comparison with vacuum energy. Embedding this structure into a four–form sequester framework, we showed that global constraints neutralize constant matter contributions inside the invariant combination $T^{(m)}_{\mu\nu}-\Lambda g_{\mu\nu}$, while the geometric unit operator does not couple to the global variable and therefore cannot be removed. The local field equations then take the renormalized effective form,
$M_{*}^{2}(\mu)\,G_{\mu\nu}=\bigl[T^{(m)}_{\mu\nu}-\Lambda g_{\mu\nu}\bigr]-T_{s}(\mu)\,g_{\mu\nu}+H_{\mu\nu}[R^{2},\nabla R,\ldots]$
with higher–derivative tensors suppressed at low curvature. Using the background field method we argued that graviton loops and mixed matter–graviton loops renormalize the unit operator, so $T_{s}(\mu)$ runs with the scale. The residual constant source may be identified with dark energy at an infrared matching scale relevant for cosmology.
An immediate next step is to extend the elastic space picture by introducing bulk modulus and bulk viscosity to space. In a homogeneous expansion a nonzero bulk viscosity produces an effective negative pressure that can evolve with the Hubble rate. This offers a natural route to time dependent dark energy and can help account for DESI indications of evolving dark energy. We will develop this viscous elastic space framework next.
7 Acknowledgment
The author acknowledges the use of ChatGPT-5 (August 2025 version) to edit and refine tone and clarity in selected sections of this manuscript. The model was used only to polish and rephrase text supplied by the author. All LLM assisted passages were reviewed and edited by the author, exercising the final editorial judgment.